### v2

# Real-World Remote Sensing Image Dehazing: Benchmark and Baseline

Abstract
Remote Sensing Image Dehazing (RSID) poses significant challenges in real-world scenarios due to the complex atmospheric conditions and severe color distortions that degrade image quality. The scarcity of real-world remote sensing hazy image pairs has compelled existing methods to rely primarily on synthetic datasets. However, these methods struggle with real-world applications due to the inherent domain gap between synthetic and real data. To address this, we introduce Real-World Remote Sensing Hazy Image Dataset (RRSHID), the first large-scale dataset featuring real-world hazy and hazy-free image pairs across diverse atmospheric conditions. Based on this, we propose MCAF-Net, a novel framework tailored for real-world RSID. Its effectiveness arises from three innovative components: Multi-branch Feature Integration Block Aggregator (MFIBA), which enables robust feature extraction through cascaded integration blocks and parallel multi-branch processing; Color-Calibrated Self-Supervised Attention Module (CSAM), which mitigates complex color distortions via self-supervised learning and attention-guided refinement; and Multi-Scale Feature Adaptive Fusion Module (MFAFM), which integrates features effectively while preserving local details and global context. Extensive experiments validate that MCAF-Net demonstrates state-of-the-art performance in real-world RSID, while maintaining competitive performance on synthetic datasets. The introduction of RRSHID and MCAF-Net sets new benchmarks for real-world RSID research, advancing practical solutions for this complex task. The code and dataset are publicly available at https://github.com/lwCVer/RRSHID.
Index Terms:
Remote sensing, real-world dehazing, color distortion, multi-branch processing, adaptive fusion.
I Introduction
Remote sensing (RS) imagery serves as a pivotal enabler for environmental sciences and agricultural applications, driving significant advancements in urban planning, agricultural monitoring, disaster management, and environmental conservation [1, 2, 3, 4, 5]. However, the ubiquitous presence of atmospheric haze presents a persistent obstacle, degrading image quality by diminishing contrast and obscuring fine details. Such degradation severely compromises the precision and reliability of subsequent analytical tasks, underscoring the urgent need for efficient, robust dehazing algorithms tailored to practical RS applications.
Figure 1: Visual comparison of synthetic and real-world RS hazy images. (a) Synthetic hazy images from the RS-HAZE dataset [6]. (b) Real-world hazy images from our RRSHID dataset, highlighting complex color variations.
In recent years, deep learning has catalyzed remarkable progress in RS image dehazing (RSID) [7]. Pioneering efforts such as TransWeather [8] introduced Transformer architectures to dehazing, while 4KDehazing [9] advanced high-resolution dehazing through sophisticated multi-scale feature extraction. These methods have predominantly been developed and validated on synthetic datasets like RICE [10] and RS-HAZE [6], achieving impressive results—e.g., FFA-Net [11] with a Peak Signal-to-Noise Ratio (PSNR) of 33.52 dB on RICE, and DehazeFormer [6] with 39.57 dB on RS-HAZE. However, when deployed on real-world RS hazy images, these models frequently suffer substantial performance drops. This gap arises from three critical challenges:
  1. 1.
    Domain Discrepancy: Synthetic datasets rely on simplified atmospheric scattering models [12], which inadequately represent the intricate, non-linear atmospheric dynamics of real-world scenes.
  2. 2.
    Complex Atmospheric Variability: Real-world RS imagery contends with altitude-dependent atmospheric effects and spatially heterogeneous haze distributions [13], which vary with elevation, time, and geography—conditions significantly more complex than those in synthetic data [6].
  3. 3.
    Color Distortion Challenges: Real-world hazy images display pronounced color distortions caused by multiple scattering and sensor-specific responses [14], as shown in Fig. 1, in stark contrast to the basic color shifts observed in synthetic datasets.
To address these challenges, we introduce the Real-World Remote Sensing Hazy Image Dataset (RRSHID). It is the first large-scale dataset featuring real-world hazy and hazy-free image pairs across diverse atmospheric conditions and geographical contexts. Its construction is detailed in Section IV.
With this dataset established, we propose the Multi-branch Color-calibrated Adaptive Fusion Network (MCAF-Net), which demonstrates state-of-the-art (SOTA) performance in real-world RSID challenges.
At its core, MCAF-Net utilizes the Multi-branch Feature Integration Block Aggregator (MFIBA) as its primary encoding component, consisting of three cascaded Multi-branch Feature Integration Blocks (MFIBs). Each MFIB splits input features into four parallel branches—height-width (hw), channel-height (ch), channel-width (cw), and point-wise integration—using a channel-splitting strategy [15]. Tailored convolution and adaptive parameters within each branch extract multi-scale, multi-dimensional feature representations, with grouped convolutions and parameter sharing ensuring computational efficiency. By cascading three MFIBs, MFIBA enhances feature robustness and regularizes information flow, adeptly balancing efficiency and representational capacity for real-world RSID.
Traditional RSID methods often adopt simplistic reconstruction techniques, such as direct feature mapping [16] or basic skip connections [17], neglecting the pivotal role of color calibration in real-world contexts. To address this, we introduce the Color-calibrated Self-supervised Attention Module (CSAM). CSAM combines self-supervised color learning with attention-driven refinement via two mechanisms: (1) an intermediate ”fake image” generated through a learnable color correction matrix for implicit color supervision, and (2) a query-key-value attention mechanism where color-calibrated features guide refinement through a global channel attention branch (for broad color consistency, e.g., sky regions) and a local depth-wise convolution branch (for detail preservation, e.g., urban edges). Tailored for RSID and inspired by recent restoration advances [18], CSAM bridges encoder and decoder, ensuring both color accuracy and structural integrity.
Moreover, conventional feature fusion in RSID—often limited to 1$\times$1 convolutions or selective kernel methods [6]—struggles with the spatial heterogeneity and scale-varying degradation of real-world haze. These approaches either oversimplify feature interactions or rely on static receptive fields, failing to capture spectral-spatial complexity. Our Multi-scale Feature Adaptive Fusion Module (MFAFM) overcomes these shortcomings by integrating features across scales and layers via a three-branch architecture (3$\times$3, 5$\times$5, 7$\times$7 kernels). Dual attention—channel and spatial—prioritizes salient features, while residual learning and adaptive adjustments stabilize training and align features effectively. This design ensures adaptive, multi-scale fusion, preserving local details and global coherence critical for real-world RSID.
Our main contributions are summarized as follows:
  1. 1.
    RRSHID as the first real-world RS hazy image dataset, highlighting how it addresses fundamental limitations in existing synthetic datasets.
  2. 2.
    MCAF-Net specifically tailored to the challenges of real-world RSID, rather than simply adapting general image dehazing techniques.
  3. 3.
    Our proposed modules (MFIBA, CSAM, and MFAFM) that specifically address the unique characteristics of real-world RS scenarios, including complex haze distributions, color distortion issues, and the need to preserve both fine details and global context.
  4. 4.
    The practical significance of our approach for real-world RS applications, emphasizing computational efficiency and performance improvements in authentic scenarios.
II Related Work
II-A Model-Based RSID
Model-based methods for RSID have historically leveraged physical models of atmospheric scattering to estimate and mitigate haze effects. Pioneering efforts, such as the Dark Channel Prior (DCP) [7], relied on the assumption that haze-free images contain dark pixels to derive transmission maps. However, this assumption often falters in RS scenarios due to expansive bright regions, such as deserts or snow-covered landscapes. To address this, subsequent innovations like the Haze-Lines Prior [19] exploited linear pixel relationships in hazy images, while the Color Attenuation Prior [20] utilized depth-color attenuation correlations to refine estimates. Despite these advancements, such methods frequently struggle to accommodate the diverse atmospheric conditions and intricate scene compositions prevalent in RS imagery. More recent developments, exemplified by Non-Local Image Dehazing [21], have integrated spatial context and global information to bolster performance. Nevertheless, model-based approaches remain challenged by varying atmospheric dynamics and the preservation of fine details in RS images.
II-B Deep Learning-Based RSID
The advent of deep learning has transformed RSID, surpassing the limitations of traditional model-based techniques. Early contributions, such as DehazeNet [22] and MSCNN [23], pioneered convolutional neural networks (CNNs) for transmission map estimation, establishing a benchmark for superior performance. AOD-Net [24] advanced this paradigm by introducing end-to-end training for clean image recovery. As the field matured, attention mechanisms and multi-scale architectures gained traction. FFA-Net [11] and GridDehazeNet [25] incorporated these features to effectively capture global context and local intricacies. RSDehazeNet [26] and HyperDehazing [27] introduced attention mechanisms specifically designed for multispectral and hyperspectral RS images. PSMB-Net [28] extended this trajectory with dual-path feature interactions to tackle non-uniform haze in urban RS scenes, while DR3DF-Net [29] enhanced feature integration through dynamic routing, enabling adaptive fusion across scales.
Recent innovations have shifted toward transformer-based architectures, marking a significant evolution in RSID. DehazeFormer [6] leveraged shifted window attention to model long-range dependencies, offering robust haze removal across complex scenes. RSDformer [30] further enhanced image structure recovery by incorporating novel self-attention mechanisms to capture both local and global correlations. PCSformer [31] built upon this by introducing cross-stripe attention modules to capture vertical and horizontal haze gradients, augmented by proxy tokens for computational efficiency. Trinity-Net [1] integrated gradient-guided Swin transformers with physics-informed priors, facilitating joint optimization of dehazing and downstream tasks, such as land cover classification. Concurrently, state space models have emerged as compelling alternatives to traditional CNNs and transformer architectures. Notably, RSDehamba [32] and HDMba [33] pioneered the application of the Mamba architecture to RSID and hyperspectral RSID, respectively, demonstrating superior efficiency in modeling long-range dependencies with linear computational complexity.
To overcome data scarcity—a persistent challenge in RSID—hybrid learning strategies have emerged. SCANet [34] proposed a self-paced semi-curricular attention network, blending curriculum learning with channel-spatial attention to address non-homogeneous haze. Semi-supervised methods [35] harnessed both labeled and unlabeled data, while domain adaptation techniques [13] bridged synthetic and real-world domains. Self-supervised learning [36] has also shown promise in exploiting unlabeled RS data for improved dehazing. Recent unpaired training approaches have effectively addressed the synthetic-real domain gap. Diff-Dehazer [37] integrates diffusion priors with CycleGAN for bijective mapping, leveraging both physical priors and cross-modal representations. DNMGDT [38] utilizes a parameter-shared architecture with adaptive weighting of prior-based pseudo labels and physical model guided domain transfer to enhance generalization to real hazy scenes. SFSNiD [39] explored frequency-consistent twin fusion, while MaxDehazeNet [40] applied maximum flow theory to model atmospheric scattering, enhancing physical interpretability.
Figure 2: The proposed MCAF-Net architecture. It adopts a U-shaped design for real-world RSID. The encoder utilizes the MFIBA with three cascaded MFIBs, while the decoder integrates CSAM and MFAFM, creating a robust framework for real-world dehazing.
III Proposed Method
In this section, we detail the architecture and key components of the proposed MCAF-Net, specifically engineered for RSID. As illustrated in Fig. 2, MCAF-Net employs a U-shaped architecture with five stages, integrating three novel modules: MFIBA, CSAM and MFAFM.
The encoder leverages MFIBA to perform efficient, cross-dimensional feature extraction through cascaded multi-branch processing, addressing the spatial heterogeneity of real-world haze. CSAM bridges the encoder and decoder by progressively refining color distortion through attention-guided refinement, ensuring structural preservation. In the decoder, MFAFM adaptively fuses multi-scale features using dual attention mechanisms—channel attention for color channel recalibration and spatial attention for terrain-specific haze suppression—effectively balancing local details and global context. This design is tailored to tackle the complex atmospheric variability and color distortions prevalent in real-world RSID scenarios, outperforming conventional approaches reliant on static feature processing.
Figure 3: Structure of the proposed MFIB. It employs axis-oriented attention to refine feature representation and flow, achieving superior dehazing performance with reduced computational overhead.
III-A MFIBA
Conventional RSID methods uniformly process features across all dimensions, propagating redundant information that obscures critical structural details and degrades generalization in complex atmospheric conditions. Real-world RSID requires robust feature extraction capable of addressing spatially heterogeneous haze distributions, a challenge unmet by conventional methods reliant on static full-channel convolutions [41, 11]. To overcome this limitation, we propose the MFIBA, a lightweight module grounded in dynamic feature disentanglement theory [42]. MFIBA selectively amplifies haze-invariant features while suppressing noise, achieving a balance between computational efficiency and representational capacity tailored for real-world RSID.
MFIBA processes input features $\mathbf{X}\in\mathbb{R}^{b\times c\times h\times w}$ by partitioning them into four subgroups $\{x_{1},x_{2},x_{3},x_{4}\}$, each with dimensions $(1,c/4,h,w)$. Three learnable queries—$\mathbf{Q}_{hw}$, $\mathbf{Q}_{ch}$ and $\mathbf{Q}_{cw}$—are dynamically generated for the spatial ($h$-$w$), channel-height ($c$-$h$), and channel-width ($c$-$w$) dimensions. All queries are independently randomly initialized. In multi-branch partial queried learning, we transform ${x_{1}}$, ${x_{2}}$, and ${x_{3}}$ into formats $(b,c,h,w)$, $(b,w,c,h)$, and $(b,h,c,w)$, respectively. We then apply bilinear interpolation to adjust the last two dimensions of queries to match feature groups, and employ depth-wise convolutions to enhance query representations. These queries interact with reshaped feature groups via Hadamard products $\mathbf{Q}_{ij}\odot\mathbf{X}_{ij}$, selectively attenuating redundant haze patterns while preserving cross-dimensional correlations. For ${x_{4}}$, we solely apply 1$\times$1 convolution for channel-wise information interaction, achieving partial self-relevant feature representation. Subsequently, the four groups of enhanced features are concatenated and transformed through a simple feed-forward network. A cascaded architecture with three MFIBs stages progressively refines contextual representations through residual connections, expanding the effective receptive field to capture large-scale haze structures. Computational efficiency is achieved through depth-wise separable convolutions [43] and 1$\times$1 channel interactions, reducing complexity from $\mathcal{O}(c^{2}hwk^{2})$ to $\mathcal{O}(c^{2}hw/4+3chw)$. This process is formalized as:
$\vspace{-1pt}\begin{split}&[X_{hw},X_{ch},X_{cw},X_{0}]=\text{LN}(\mathbf{X})\\ &\mathbf{Q}_{hw},\mathbf{Q}_{ch},\mathbf{Q}_{cw}=\text{DWConv}_{3\times 3}(\mathbf{Q})\\ &\mathbf{X}^{\prime}_{queried}=\text{Concat}(\mathbf{Q}_{ij}\odot\mathbf{X}_{ij},\text{Conv}_{1\times 1}(X_{0}))\\ &\mathbf{X}_{queried}=\text{MLP}(\mathbf{X}^{\prime}_{queried})+\mathbf{X}\end{split}\vspace{-1pt}$ (1)
where $\text{LN}(\cdot)$ denotes layer normalization, $\text{DWConv}(\cdot)$ is depth-wise convolution and $i,j\in\{hw,ch,cw\}$. The MFIBA further enhances contextual information by cascading multiple MFIBs. Its structure as follows:
$\vspace{-3pt}\text{MFIBA}(\mathbf{X})=\text{Conv}_{3\times 3}(\text{MFIB}_{3}(\text{MFIB}_{2}(\text{MFIB}_{1}(\mathbf{X}))))$ (2)
By synergizing dynamic dimension-wise attention, multi-scale feature fusion, and lightweight operators, MFIBA establishes a theoretically grounded framework for efficient and adaptive dehazing, effectively addressing the spatiotemporal heterogeneity of real-world RS imagery without compromising computational practicality.
III-B CSAM
RS images under hazy conditions often exhibit severe color distortion, which is more complex compared to near-ground images, as atmospheric thickness and composition vary significantly at different altitudes. Conventional RSID [44, 45, 46] methods often overlook color calibration during early decoding stages [47], limiting their effectiveness against the severe color distortions [14] induced by atmospheric scattering in real-world scenarios. To address this, we introduce CSAM, integrated across multiple decoder levels, to progressively refine haze-free features while ensuring color accuracy and structural integrity, as shown in Fig. 4.
CSAM enables progressive refinement of haze-free features through three key operations. First, PixelShuffle upsampling reconstructs high-resolution features, preserving channel-wise relationships. Second, a learnable color correction matrix $W_{\text{color}}\in\mathbb{R}^{3\times C}$ transforms features via $I_{\text{fake}}=f_{\text{color}}(\text{Conv}_{1\times 1}(X_{\text{up}}))$, where $f_{\text{color}}(\cdot)$ implements spectral compensation through matrix multiplication and bias addition. This operation simulates the inverse process of atmospheric scattering by adaptively adjusting RGB channel gains and offsets [48]. Third, the attention employs query-key-value (QKV) processing with depth-wise convolution refinement:
$\begin{split}\vspace{-3mm}X_{up}=U(X)&,\hskip 11.38109ptQ,K=\text{Conv}_{1\times 1}(\text{LN}(X_{up}))\\ A&=\text{softmax}(\alpha\cdot\frac{QK^{T}}{\sqrt{d_{k}}})\\ F_{c}&=AV+\text{DWConv}_{3\times 3}(V)\end{split}$ (3)
Here, $U(\cdot)$ denotes the upsampling operation. $\alpha$ is a learnable parameter and $d_{k}$ is the dimension of the key vectors. $V$ is the value vector and $\text{DWConv}_{3\times 3}$ is a depth-wise separable convolution which preserves spatial details through parameter-efficient filtering, while the attention map $A\in\mathbb{R}^{H\times W\times C}$ emphasizes haze-invariant channels (e.g., near-infrared bands less affected by scattering). This dual-stream design is biologically inspired by the parvocellular pathway for color processing and magnocellular pathway for spatial acuity in human vision [49].
Figure 4: Architecture of CSAM. The color calibration branch (top) learns sensor-specific spectral adjustments via $W_{\text{color}}$, while the attention branch (bottom) integrates global channel dependencies and local spatial details via QKV processing.
Our proposed CSAM module offers two significant contributions. First, it operates specifically in the color space with a learnable correction matrix that adaptively adjusts color distribution based on input characteristics, addressing complex color distortion in haze-affected RS imagery. Second, CSAM implements an innovative self-attention mechanism that differs from similar strategies optimized for object tracking[50, 51] by: (1) focusing on color rather than spatial features to address differential atmospheric scattering effects across color channels; (2) maintaining multi-scale color consistency through MFIBA module integration; and (3) specifically targeting real-world RS dehazing challenges rather than general vision tasks. These design elements enable CSAM to effectively address color restoration challenges in RSID, achieving superior color correction while preserving critical image details.
III-C MFAFM
Traditional RSID feature fusion methods, such as 1$\times$1 convolutions and SKFusion [6], struggle to handle altitude-dependent haze variations—where haze density increases exponentially at lower altitudes —and fail to adaptively balance fine details (e.g., urban edges) with the global atmospheric context. Fixed-scale operators cannot jointly address these spatially heterogeneous phenomena, leading to over-smoothing or residual artifacts. The proposed MFAFM module addresses the critical challenge of multi-scale feature fusion in real-world RSID, where altitude-dependent haze density variations and complex land cover patterns demand adaptive integration of spatial-contextual information across scales. As shown in Fig. 5, the module processes three inputs: the current feature map $x$ and two skip connections $skip1$, $skip2$ from preceding layers, each capturing distinct scale-specific characteristics.
MFAFM employs convolutional kernels of sizes 3$\times$3, 5$\times$5, and 7$\times$7 to capture scale-specific haze patterns: small kernels preserve high-frequency details (e.g., building edges), medium kernels address mid-range gradients (e.g., vegetation clusters), and large kernels aggregate global context (e.g., sky-to-ground transitions). These operations are defined as:
$\begin{split}x_{1}&=\text{Conv}_{3\times 3}(x)\\ x_{2}&=\text{Conv}_{5\times 5}(\text{$skip1$})\\ x_{3}&=\text{Conv}_{7\times 7}(\text{$skip2$})\end{split}$ (4)
The resulting feature maps are then concatenated. To enhance channel-wise feature relationships, we employ a channel attention mechanism $\text{CA}(\cdot)$ that recalibrates RGB channel relationships to address color distortion. Spatial relationships, which identifies and processes regions with varying haze densities across different terrain types, are further refined using a spatial attention mechanism $\text{SA}(\cdot)$:
$\begin{split}\ \text{X}_{\text{cat}}&=\text{Concat}(x_{1},x_{2},x_{3})\\ \text{X}_{\text{ca}}&=\text{X}\odot\text{CA}(\text{X}_{\text{cat}})\\ \text{X}_{\text{sa}}&=\text{X}_{\text{ca}}\odot\text{SA}(\text{X}_{\text{ca}})\end{split}$ (5)
The refined features are then fused with the original input through residual connections. MFAFM can be summarized as:
$\text{MFAFM}(x,\text{$skip1$},\text{$skip2$})=\text{Conv}_{1\times 1}(\text{X}_{\text{sa}})+\text{Conv}_{1\times 1}(x)$ (6)
ensuring stable gradient flow while preserving critical high-resolution details. This hierarchical fusion strategy emulates the human visual system’s multi-scale processing mechanism, where retinal ganglion cells progressively integrate local contrast and global luminance information.
Figure 5: Architecture of MFAFM. Three convolutional branches capture scale-specific features, while dual attention mechanisms suppress haze-corrupted components. Residual connections preserve original feature integrity.
III-D Loss Function
MCAF-Net is optimized using a hybrid loss function combining $L_{2}$ and perceptual losses. While $L_{2}$ loss aligns predictions with ground truth for high PSNR, it may weaken structural similarity (SSIM). The perceptual loss, extracted from the 3rd, 8th, and 15th layers of a pretrained VGG16 network, enhances structural similarity and visual quality:
$L_{\text{perceptual}}(\tilde{J},J)=\sum_{l\in\{3,8,15\}}L_{2}(\text{VGG}_{l}(\tilde{J}),\text{VGG}_{l}(J))$ (7)
Perceptual loss aids in preserving the structural and detail information of the images, thereby improving the visual quality of the dehazed images. The final loss function combines these two losses, expressed as:
$L(\tilde{J},J)=L_{2}(\tilde{J},J)+\lambda L_{\text{perceptual}}(\tilde{J},J)$ (8)
where $\lambda=0.04$ balances the two components [8, 52].
IV RRSHID: A Real-World RSID Dataset
RSID is a persistent challenge, particularly in real-world contexts where atmospheric conditions exhibit significant complexity and variability. While deep learning has driven notable progress in this domain, its application to authentic RS imagery remains constrained. Studies by Li et al. [41] and Qin et al. [11] demonstrate that existing methods often falter under diverse atmospheric scattering patterns and heterogeneous scene compositions. Furthermore, Zhang et al. [53] underscore the limited generalization of many models to real-world RS data, attributing this to the reliance on synthetic training datasets that fail to capture the intricacies of natural haze. To bridge this critical gap, we present the Real-World Remote Sensing Hazy Image Dataset (RRSHID), a pioneering collection of 3,053 paired hazy and haze-free images. spanning urban, agricultural, and coastal landscapes across China, including provinces such as Anhui, Sichuan, Shanghai, Shandong, and Fujian, these images were captured under diverse illumination and seasonal conditions. Unlike synthetic datasets such as those in [10, 6, 54, 1], which depend on simplified atmospheric models, RRSHID leverages collaborative data acquisition with meteorological agencies. RRSHID ensures realistic representations of altitude-dependent haze distributions and sensor-induced color distortions. It serves as a vital benchmark for evaluating dehazing algorithms and advancing domain generalization in real-world RSID.
IV-A Image Acquisition and Processing
The development of RRSHID involved a systematic methodology for image selection and preprocessing, implemented through collaboration with remote sensing domain experts. Working in concert with meteorological agencies and geospatial scientists, we employed multi-temporal satellite image alignment techniques to identify spatially congruent regions captured by satellite platforms across multiple acquisition cycles. To ensure the authenticity and precision of our paired images, we developed a comprehensive dataset generation framework consisting of three key stages:
IV-A1 Data Source Selection
The dataset draws from the GF_PMS: L1A optical satellite, part of the Optical and SAR Satellite Payload Retrieval system, which delivers high-resolution imagery at 1-meter panchromatic and 4-meter multispectral scales. Covering diverse terrains across China—urban centers, agricultural fields, and coastal zones—the images were acquired 2021-2023, reflecting seasonal and illumination variability. Each raw satellite image, stored in GeoTIFF format with an approximate file size of 8 GB, contains rich metadata including acquisition time, sensor parameters, and precise geolocation information.
IV-A2 Temporal-Spatial Alignment
The critical challenge in creating paired hazy/haze-free images lies in ensuring precise alignment between degraded and clear images. Our approach addresses this by a dual temporal-spatial alignment strategy:
  • Temporal Proximity Selection: We first identified temporally proximate multispectral image pairs (typically within 1-3 months) from the same geographic region that exhibited contrasting atmospheric conditions—one with significant haze presence and one with clear visibility. This temporal proximity minimizes seasonal variations in vegetation, building structures, and other land features.
  • Geospatial Coordinate Matching: For each identified image pair, we implemented precise geospatial alignment by extracting geographic coordinate metadata (latitude/longitude) from image headers and identifying overlapping regions with consistent land features but different atmospheric conditions. Through precise marking of the top-left and bottom-right geographic coordinates of target regions, we ensured exact spatial correspondence before applying coordinate-based cropping to extract identical geographic areas from both images.
IV-A3 Image Preprocessing
Following alignment, each image pair underwent a standardized preprocessing workflow:
  • Channel Conversion: Multispectral images with multiple bands were transformed into 3-channel RGB (red, green, blue) format using Adobe Photoshop, prioritizing bands optimal for visual analysis.
  • Format Optimization: Custom Python scripts refined the cropped images, generating $256\times 256$ pixels subimages and converting them from TIFF to PNG format to enhance compatibility with machine learning frameworks.
TABLE I: Comparison of RRSHID with Existing Remote Sensing Datasets
Dataset Size Real/Synthetic Paired Resolution Application Color Bias Image Type Scene Diversity
StateHaze1K [54] 1,200 Synthetic Yes 512$\times$512 Dehazing No RGB Single
RSID [1] 1,000 Synthetic Yes 256$\times$256 Dehazing No RGB Single
RS-HAZE [6] 51,300 Synthetic Yes 512$\times$512 Dehazing No RGB Single
LHID [55] 31,017 Synthetic Yes 512$\times$512 Dehazing No RGB Multiple
DHID [55] 14,990 Synthetic Yes 512$\times$512 Dehazing No RGB Single
HyperDehazing [27] 2,000 Synthetic Yes 512$\times$512 Dehazing No Hyperspectral Multiple
HyperDehazing [27] 70 Real No 512$\times$512 Dehazing No Hyperspectral Multiple
RICE1 [10] 500 Real Yes 512$\times$512 Dehazing No RGB Single
RICE2 [10] 736 Real Yes 512$\times$512 Cloud/Haze Removal No RGB Single
HRSI [56] 796 Real No 512$\times$512 to 4000$\times$4000 Dehazing (Object Detection) No RGB Single
RRSHID 3,053 Real Yes 256$\times$256 Dehazing Yes RGB Multiple
Figure 6: Haze density distribution across the RRSHID dataset.
IV-B Haze Density Estimation
To classify images by haze intensity, we adopted the dark channel prior method introduced by He et al. [7]. This technique estimates the transmission map $t(x)$ based on the principle that, in haze-free images, most non-sky regions exhibit at least one color channel with near-zero intensity. The dark channel is mathematically defined as:
$J_{\text{dark}}(x)=\min_{y\in\Omega(x)}\left(\min_{c\in\{r,g,b\}}J^{c}(y)\right),$ (9)
where $J^{c}$ denotes a color channel of the haze-free image $J$, and $\Omega(x)$ represents a local patch centered at pixel $x$. The transmission map is subsequently derived as:
$\tilde{t}(x)=1-\omega\min_{y\in\Omega(x)}\left(\min_{c}\frac{I^{c}(y)}{A^{c}}\right),$ (10)
where $I$ is the hazy image, $A$ is the atmospheric light, and $\omega$ is a constant (typically 0.95) to retain subtle haze effects. For each image, we calculated the mean dark channel value as a proxy for haze density. The threshold determination for haze categorization followed a data-driven approach. We performed k-means clustering (k=3) on the distribution of mean dark channel values across the entire dataset, which revealed natural groupings at 110.58 and 159.31. These thresholds were further validated through visual assessment by remote sensing experts and correlation with meteorological visibility data from ground stations (Pearson’s r = 0.87), as depicted in Fig. 6. Images were accordingly categorized into thin (below 110.58), moderate (110.58–159.31), and thick (above 159.31) haze levels.
IV-C Dataset Characteristics
TABLE II: Distribution of Image Pairs in the RRSHID Dataset
Haze Level Total Train Test Validation
Thin haze 763 610 76 77
Moderate haze 1,526 1,220 152 154
Thick haze 764 611 76 77
Total 3,053 2,441 304 308
RRSHID encompasses 3,053 pairs of hazy and haze-free RS images, constituting the largest real-world dataset of its kind, as comparison in Table I. These pairs are stratified by haze density into three categories, as detailed in Table II.
Each image, standardized at $256\times 256$ pixels, strikes a balance between spatial detail and computational efficiency for deep learning applications. As the first large-scale, real-world paired RS dataset, RRSHID captures authentic atmospheric phenomena absent in synthetic counterparts, including:
· Heterogeneous haze densities and spatial distributions within individual images,
· Intricate interactions between haze and diverse land cover types (e.g., urban, vegetative, and marine),
· Color deviations caused by variations in natural lighting conditions and atmospheric composition.
These properties, as depicted in Fig. 7, render RRSHID a crucial resource for developing robust RSID algorithms, assessing haze impacts on applications such as object detection [57, 58] and change detection [59, 60], and deepening insights into atmospheric effects on satellite imagery.
Figure 7: Sample of the RRSHID that we proposed. Unlike synthetic datasets (e.g., RSID, RS-HAZE), RRSHID demonstrates authentic characteristics in haze density distribution, land cover diversity, and color distortion.
V Experiments
V-A Experimental Settings
V-A1 Datasets
MCAF-Net is evaluated on the RRSHID dataset and five publicly available benchmarks: RSID [1], RICE1 [10], RICE2 [10], StateHaze1K-thick [54], and HRSI100. These datasets are selected for their diverse characteristics:
RSID: Includes 1000 hazy and haze-free image pairs, with 900 used for training and 100 for testing.
RICE1: Comprises 500 cloudy and cloud-free pairs, split into 400 for training and 100 for testing.
RICE2: Features 736 pairs with complex thick clouds and shadows, divided into 588 for training and 147 for testing.
StateHaze1K-thick: A subset of 400 hazy pairs, allocated as 320 for training, 45 for testing, and 35 for validation.
HRSI100: Contains 100 real-world RS hazy images cropped to $256\times 256$ resolution from the HRSI [56] dataset without ground truth, specifically designed to evaluate generalization capability in real-world scenarios.
The RRSHID dataset, designed to reflect real-world RS challenges, complements these synthetic and real-world benchmarks to comprehensively assess generalization performance.
V-A2 Implementation Details
All models trained for 300 epochs on a single NVIDIA Tesla P100-PCIE GPU. We configure the patch size to $256\times 256$ and the batch size to 8. The Adam optimizer is utilized with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, paired with an initial learning rate of 0.001, which decays to $1\times 10^{-8}$ via a cosine annealing scheduler. The network architecture comprises five stages, with embedding dimensions of [24, 48, 96, 48, 24] and corresponding depths of [8, 8, 16, 8, 8], optimizing both capacity and efficiency for RSID.
V-A3 Evaluation Metrics
To assess the performance on supervised hazy datasets, we employed four metrics: PSNR, Structural Similarity Index Measure (SSIM), Mean Squared Error (MSE), and Learned Perceptual Image Patch Similarity (LPIPS). Higher PSNR and SSIM values signify improved image quality, while lower MSE and LPIPS indicates enhanced dehazing efficacy. To further validate generalization capability in real-world scenarios, we additionally utilized Natural Image Quality Evaluator (NIQE) [61] and Integrated Local NIQE (IL-NIQE) [62] metrics, where lower values indicate better perceptual quality without requiring reference images.
V-B Experimental Results
TABLE III: Performance comparison of different methods on RRSHID dataset across various haze levels. Values in red and blue denote the best and second-best performance, respectively.
Method RRSHID-thin RRSHID-moderate RRSHID-thick RRSHID-average
PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS
DCP[7] 18.46 0.4564 0.0192 0.4851 17.80 0.4856 0.0238 0.4700 18.39 0.4843 0.0208 0.4996 18.22 0.4754 0.0213 0.4849
FFA-Net [11] 17.08 0.4452 0.0326 0.5761 17.40 0.5385 0.0346 0.5450 16.71 0.4792 0.0377 0.5573 17.06 0.4876 0.0350 0.5595
GridDehazeNet[25] 22.77 0.6145 0.0069 0.4123 22.62 0.6468 0.0083 0.3833 23.96 0.7112 0.0061 0.3947 23.12 0.6575 0.0071 0.3968
4KDehazing[9] 22.83 0.6177 0.0063 0.4352 22.47 0.6505 0.0083 0.4590 22.55 0.6912 0.0099 0.4754 22.62 0.6531 0.0082 0.4565
SCAnet[34] 18.37 0.4718 0.0200 0.4827 18.11 0.0538 0.0210 0.4962 19.07 0.5966 0.0180 0.4734 18.52 0.3741 0.0195 0.4841
Trinity-Net[1] 20.51 0.5728 0.0120 0.4578 22.46 0.5728 0.0085 0.4314 24.11 0.7234 0.0058 0.4206 22.36 0.6230 0.0060 0.4366
DehazeFormer[6] 22.74 0.6005 0.0071 0.4609 23.06 0.6137 0.0076 0.4495 24.69 0.7143 0.0051 0.4377 23.50 0.6428 0.0066 0.4494
PCSformer[31] 21.83 0.5427 0.0088 0.4963 22.09 0.5984 0.0092 0.5021 23.71 0.6547 0.0055 0.4996 22.54 0.5996 0.0070 0.4993
PhDnet[63] 22.64 0.6054 0.0072 0.4658 22.92 0.6448 0.0075 0.4019 24.28 0.6996 0.0053 0.3993 23.28 0.6499 0.0067 0.4223
MCAF-Net 23.32 0.6236 0.0059 0.4023 23.60 0.6583 0.0063 0.3799 25.40 0.7221 0.0040 0.3942 24.11 0.6680 0.0054 0.3921
Figure 8: Visual samples generated by the evaluated methods on the RRSHID dataset. Note that beneath each hazy or restored image, we present the corresponding error maps depicting discrepancies with the ground truth.
V-B1 Real-World RSID
To evaluate MCAF-Net’s effectiveness in real-world RSID, we conducted comprehensive experiments on our RRSHID dataset. This dataset, specifically curated to represent challenging real-world scenarios, served as a robust benchmark for assessing dehazing algorithms in RS imagery.
Table III presents a detailed performance comparison of the proposed MCAF-Net against several SOTA dehazing methods on the RRSHID dataset, which encompasses real-world RS images under varying haze conditions: thin, moderate, and thick. The results, as shown in the table, indicate that MCAF-Net consistently achieves the highest performance across all haze levels and metrics, marked in red, with the second-best performances highlighted in blue.
RRSHID-thin: MCAF-Net achieves a PSNR of 23.32, an SSIM of 0.6236, an MSE of 0.0059 and a LPIPS of 0.4023, surpassing all competing methods. The second-best performer, 4KDehazing [9], records a PSNR of 22.83 and an SSIM of 0.6177, trailing MCAF-Net by 0.49 in PSNR and 0.0059 in SSIM, with a higher MSE of 0.0063 and LPIPS of 0.4352.
RRSHID-moderate: MCAF-Net excels with a PSNR of 23.60, an SSIM of 0.6583, an MSE of 0.0063 and a LPIPS of 0.3799. In comparison, DehazeFormer [6], the second-best method, achieves a PSNR of 23.06 and an SSIM of 0.6137, falling short by 0.54 in PSNR and 0.0446 in SSIM, with an MSE of 0.0076 and a LPIPS of 0.4495.
RRSHID-thick: MCAF-Net demonstrates superior performance with a PSNR of 25.40, an SSIM of 0.7221, an MSE of 0.0040 and a LPIPS of 0.3942. DehazeFormer obtains a PSNR of 24.69 and an SSIM of 0.7143, lagging by 0.71 in PSNR and 0.0078 in SSIM, with an MSE of 0.0051 and a LPIPS of 0.4377.
RRSHID-average: Across all haze levels, MCAF-Net maintains its lead with an average PSNR of 24.11, an SSIM of 0.6680, an MSE of 0.0054 and a LPIPS of 0.3921. The second-best results vary by metric: DehazeFormer achieves a PSNR of 23.50 (0.61 lower than MCAF-Net), GridDehazeNet [25] records an SSIM of 0.6575 (0.0105 lower) and a LPIPS of 0.3968 (0.0047 higher), and Trinity-Net [1] yields an MSE of 0.0060 (0.0006 higher).
The consistent superiority of MCAF-Net across all haze conditions and evaluation metrics underscores its robustness and effectiveness in real-world RSID. Notably, the performance gap is most pronounced under thick haze conditions, where MCAF-Net achieves significant improvements in PSNR (0.71 higher than DehazeFormer) and MSE (0.0011 lower), indicating its capability to recover fine details and structural integrity in heavily degraded images.
Compared to established methods such as DCP [7], FFA-Net [11], SCAnet [34] and PCSformer [31], MCAF-Net demonstrates substantial improvements, especially in SSIM and MSE, reflecting its enhanced ability to preserve structural fidelity and minimize reconstruction errors. Even against more recent approaches such as GridDehazeNet, 4KDehazing, Trinity-Net, DehazeFormer and PhDnet[63], MCAF-Net stands out as a superior solution, reinforcing its position as a SOTA method for real-world RSID.
To provide a more comprehensive assessment beyond numerical metrics, we included visual comparisons of the dehazing results in Fig. 8. These qualitative examples offered insight into the perceptual improvements achieved by our method. Qualitative results further validate the quantitative findings, showcasing our algorithm’s ability to effectively remove haze while preserving crucial image details and enhancing overall visual clarity, while FFA-Net and 4KDehazing often leave residual haze or introduce artifacts. Despite MCAF-Net’s overall effectiveness across various haze conditions, it still faces challenges in extremely dense haze scenarios. When visibility is severely compromised, the model struggles to recover fine-grained details and occasionally produces artifacts in regions with complex textures. As demonstrated in the last row of Fig. 8, our method shows performance limitations when dealing with intricate structures such as dense urban areas or complex vegetation patterns.
In summary, the experimental results on the RRSHID dataset validate the efficacy of MCAF-Net for real-world RSID. Its consistent outperformance across diverse haze levels and metrics establishes MCAF-Net as a leading approach in this domain, offering a robust and reliable solution for enhancing the quality of haze-degraded RS imagery.
V-B2 Synthetic RSID
TABLE IV: Performance comparison of different methods on multiple haze datasets. Values in red and blue denote the best and second-best performance, respectively.
Method RSID RICE1 RICE2 Statehaze1k-thick
PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS
DCP[7] 13.87 0.6892 0.0538 0.2615 15.68 0.6860 0.0491 0.3766 14.10 0.3898 0.0656 0.6234 17.87 0.8481 0.0169 0.1879
FFA-Net[11] 18.31 0.8582 0.0204 0.1449 23.73 0.9068 0.0096 0.1885 17.77 0.6261 0.0323 0.5378 19.45 0.9023 0.0117 0.2357
GridDehazeNet[25] 23.50 0.9383 0.0080 0.0541 33.45 0.9766 0.0010 0.0405 31.54 0.8839 0.0014 0.1942 20.51 0.9097 0.0089 0.1361
4KDehazing[9] 23.61 0.9415 0.0053 0.0693 27.54 0.9425 0.0024 0.0967 25.21 0.8604 0.0087 0.3704 20.75 0.7696 0.0085 0.1674
Trinity-Net[1] 23.60 0.9322 0.0060 0.0692 23.46 0.8796 0.0100 0.0512 18.81 0.7781 0.0200 0.2287 20.43 0.8056 0.0086 0.1764
PSMB-Net[28] 25.64 0.9447 0.0030 0.0537 31.32 0.9444 0.0006 0.0428 30.21 0.8756 0.0010 0.1983 21.55 0.8489 0.0084 0.1642
DehazeFormer[6] 25.04 0.9393 0.0040 0.0557 36.15 0.9794 0.0005 0.0460 34.54 0.8875 0.0010 0.2035 22.02 0.9306 0.0063 0.1707
PCSformer[31] 23.03 0.9159 0.0063 0.0718 35.17 0.9564 0.0006 0.0421 33.80 0.8781 0.0011 0.2418 20.20 0.8125 0.0097 0.1734
PhDnet[63] 25.15 0.9406 0.0040 0.0612 35.50 0.9578 0.0005 0.0464 33.44 0.8877 0.0011 0.2277 21.99 0.8294 0.0064 0.1721
MCAF-Net 25.89 0.9531 0.0032 0.0421 35.79 0.9814 0.0005 0.0383 34.59 0.8948 0.0009 0.1964 22.16 0.9308 0.0062 0.1453
Figure 9: Visual samples generated by the evaluated methods on RSID [1], RICE1[10], RICE2[10], and Statehaze1k-thick[54] datasets.
Our method demonstrated exceptional performance in real-world RSID. To further evaluate its generalization capability, we applied the proposed approach to several publicly available synthetic RSID datasets. The methods include DCP [7], FFA-Net [11], GridDehazeNet [25], 4KDehazing [9], Trinity-Net [1], PSMB-Net [28], DehazeFormer [6], PCSformer [31], PhDnet[63] and our proposed MCAF-Net.
Theoretically, the synthetic data exhibited certain scene differences compared to the real data, which could potentially lead to suboptimal results on some public synthetic RS datasets. However, our experimental results were highly encouraging. Our method not only performed admirably on these synthetic datasets but also provided strong evidence of its robust generalization ability and resilience.
We conducted comprehensive comparative analyses across multiple SOTA methods and datasets, including RSID [1], RICE1 [10], RICE2 [10], and state1k-thick [54].
Experimental results, presented in Table IV, clearly demonstrated that MCAF-Net outperformed nearly all evaluated approaches in terms of quantitative performance on synthetic RS hazy degradation datasets. Notably, our method exhibited particularly high performance in terms of SSIM scores, indicating its superiority in preserving structural information.
Visual assessments of our method alongside other evaluated approaches were illustrated in Fig. 9, further confirming the superiority of our model. These visual comparisons highlighted our method’s ability to effectively remove haze while preserving fine details and enhancing overall image quality.
The consistent superior performance across both real-world and synthetic datasets reinforced the robustness and versatility of our proposed approach, positioning it as a promising solution for a wide range of RSID scenarios.
V-B3 The generalization verification in Real-World RSID
Figure 10: Visual samples generated by the evaluated methods on the HRSI100 dataset. Our MCAF-Net successfully eliminates haze and better restores image details, outperforming other evaluated methods.
TABLE V: Generalization comparison of state-of-the-art methods
        Method         NIQE         IL-NIQE
        DCP [7]         6.618         65.79
        FFA-Net [11]         18.609         120.47
        4KDehazing [9]         9.207         91.79
        DehazeFormer [6]         16.282         131.61
        PCSformer [31]         6.105         85.04
        PhDnet [63]         9.684         97.12
        MCAF-Net         5.607         58.32
To further validate the generalization capability of our proposed method and dataset in real-world scenarios, we conducted additional experiments on the HRSI100 dataset. For fair comparison, all methods were evaluated using models trained on the RRSHID-thick dataset without any fine-tuning. Table  V presents the quantitative results, showing that our method achieves the best NIQE and IL-NIQE scores (5.607 and 58.32 respectively), significantly outperforming other comparison methods. This demonstrates superior generalization ability on unseen real-world hazy scenes. Fig. 10 provides visual comparisons that further support our findings - our method not only effectively removes haze but also better preserves image details. These results confirm the effectiveness and generalization capability of our proposed method and dataset in real-world RS applications, which is crucial for practical deployment.
TABLE VI: Ablation study of different components.
Model PSNR SSIM MSE #Param FLOPs
Baseline 23.20 0.6952 0.0047 1.202M 51.28G
+MFIBA 24.15 0.7091 0.0053 558.1K 19.82G
+CSAM 24.61 0.7119 0.0045 582.3K 20.15G
+MFAFM 25.40 0.7221 0.0040 558.1K 19.82G
V-C Ablation Studies
The effectiveness of the proposed modules is assessed through ablation studies conducted on the RRSHID-thick dataset. These studies serve to analyze the individual contributions of each component within our proposed MCAF-Net and enable comparisons with their similar designs.
V-C1 Individual Components
In this study, we employed a baseline architecture with a U-shaped design, utilizing ResNet-Block (RNB) [64] as the basic learning block. Table  VI depicts the ablation on individual components. Initially, we replaced the RNB with our MFIBA, resulting in a performance gain of 0.95 dB in PSNR. Remarkably, this improvement was achieved with only 46.43% #Param and 38.65% FLOPs of the baseline model, demonstrating the efficiency and effectiveness of our MFIBA. Subsequently, incorporating our proposed CSAM module, the model achieved an additional performance gain of 0.46 dB, albeit with a slight increase in computational burden (582.3K #Param and 20.15G FLOPs). This result indicated the capability of our CSAM in improving performance and demonstrated the significance of color correction in RSID. Finally, we integrated our MFAFM, which led to a substantial improvement of 0.79 dB in PSNR while maintaining similar computational efficiency as MFIBA (558.1K #Param and 19.82G FLOPs). In conclusion, the ablation studies on individual components demonstrated the significant impact of each proposed component in a positive manner.
TABLE VII: Comparison of different basic learning blocks.
Component PSNR SSIM MSE #Param FLOPs
RNB [64] 24.82 0.7145 0.0048 1.15M 48.56G
FNB [65] 24.95 0.7168 0.0046 892.3K 35.82G
MFIB×$\times$1 25.13 0.7196 0.0043 508.6K 18.83G
MFIB×$\times$2 25.15 0.7207 0.0044 533.3K 19.36G
MFIB×$\times$3 25.40 0.7221 0.0040 558.1K 19.82G
MFIB×$\times$4 25.19 0.7209 0.0044 571.3K 20.02G
TABLE VIII: Performance Comparison of Different Components
Component PSNR SSIM CIEDE2000 #Param FLOPs
Baseline 23.96 0.7112 10.38 534.1K 18.50G
+ Attention 24.11 0.7234 7.37 542.3K 18.75G
+ Color Calibration 24.69 0.7143 6.27 538.5K 18.62G
+ CSAM 25.40 0.7221 5.72 558.1K 19.82G
V-C2 Basic Learning Blocks
To assess the effectiveness of our proposed MFIBA, we conducted studies comparing its performance with two commonly used learning blocks, namely RNB [64] and the FasterNetBlock (FNB) [65], in the context of RSID. Table VII presented the results of our experiments. Notably, our MFIB×$\times$3 outperformed RNB by a margin of 0.58 dB and FNB by a margin of 0.45 dB in terms of PSNR. Furthermore, we investigated the impact of cascading multiple MFIBs. As shown in Table VII, the performance gradually improved as the number of MFIBs increases, with MFIB×$\times$3 achieving the best performance while maintaining relatively low computational cost (558.1K #Param and 19.82G FLOPs). These findings demonstrated the effectiveness of our MFIBA design and its potential for enhancing RSID performance.
Figure 11: Visual comparison of CSAM components: (a)+Attention, (b)+Color Calibration, (c)+CSAM, (d) Ground Truth.
V-C3 Capacity of CSAM
To comprehensively evaluate the effectiveness of the CSAM, we conducted detailed ablation studies focusing on its two core components: the self-supervised color correction mechanism and the attention-guided feature refinement, with the results summarized in Table VIII.
The CIEDE2000 [66] metric, which quantifies perceptual color differences by considering luminance, chroma, and hue variations in the CIELAB color space, demonstrates progressive improvement through our component-wise analysis. When introducing the attention mechanism alone, we observe a moderate PSNR gain of 0.15 dB and SSIM improvement of 0.0122, primarily enhancing structural preservation. However, the substantial CIEDE2000 reduction from 10.38 to 7.37 (-29.0%) reveals remaining color distortion issues. The color calibration module shows complementary characteristics - it achieves the most significant CIEDE2000 improvement (6.27, -39.6% from baseline) through learnable color matrix transformation, though with slightly reduced SSIM performance due to enhanced color contrast. Our complete CSAM integration synergistically combines both advantages, delivering SOTA performance with 25.40 dB PSNR (+1.44 dB) and 5.72 CIEDE2000 (-45.0%), while maintaining practical complexity (558.1K parameters, +4.5% from baseline).
Fig 11 visually validates these findings. The attention-enhanced result (a) preserves building edges but shows yellowish haze residuals. The color-calibrated output (b) exhibits accurate sky color recovery at the expense of vegetation detail. Our full CSAM implementation (c) successfully integrates both capabilities, producing haze-free results with natural chromaticity comparable to GT (d).
These demonstrate CSAM’s unique capacity to bridge the gap between synthetic training and real-world deployment in RSID tasks through physics-guided color correction and data-driven feature refinement.
TABLE IX: Performance comparison of different fusion modules
Component PSNR SSIM MSE #Param FLOPs
$1\times 1$ Conv 25.01 0.7128 0.0045 563.1k 20.02G
SKFusion [6] 25.19 0.7132 0.0043 548.5k 19.42G
MFAFM 25.40 0.7221 0.0040 558.1K 19.82G
V-C4 Adaptive Feature Fusion
To comprehensively validate the efficacy of MFAFM, we performed systematic comparisons with two representative fusion approaches: standard 1$\times$1 convolution and SKFusion [6]. As quantitatively demonstrated in Table IX, the proposed MFAFM achieves marked superiority over conventional convolution operations, achieving a significant gain in PSNR of 0.39 dB. Although SKFusion marginally outperforms baseline convolution (0.12 dB improvement), our method established a new SOTA by further increasing the performance ceiling with an additional 0.21 dB enhancement. This progression not only corroborated the findings in [6] but more importantly underscored MFAFM’s exceptional capability in addressing real-world RSID challenges, confirming MFAFM’s effectiveness as a multiscale adaptive paradigm for feature fusion in RSID applications.
V-C5 Loss Functions
To investigate the effectiveness of different loss functions, we conducted ablation studies with various loss combinations. Table X presented the quantitative results. Using only $\mathcal{L}_{2}$ loss achieved a PSNR of 24.95 dB. When replacing it with perceptual loss, the PSNR decreased to 22.85 dB, but showed better perceptual quality with a SSIM of 0.6873. By combining $\mathcal{L}_{2}$ and $\mathcal{L}_{perceptual}$ losses, our model achieved the best performance with a PSNR of 25.40 dB and SSIM of 0.7221, demonstrating the complementary effects of these two loss terms in optimizing both pixel-level accuracy and perceptual quality.
TABLE X: Performance comparison of different loss functions.
Loss Function PSNR SSIM MSE
$\mathcal{L}_{2}$ 24.95 0.7217 0.0044
$\mathcal{L}_{perceptual}$ 22.85 0.6873 0.0052
$\mathcal{L}_{2}+\lambda\mathcal{L}_{perceptual}$ 25.40 0.7221 0.0040
V-D Efficiency Analysis
We conducted a comprehensive comparison of SOTA dehazing methods in Table XI, evaluating their restoration quality and computational efficiency. MCAF-Net established new benchmarks for RSID by simultaneously achieving SOTA restoration quality (PSNR 25.40 dB, SSIM 0.7221) and unparalleled efficiency (558.1K parameters, 19.82G FLOPs). While Trinity-Net processes images marginally faster (0.0400s vs. 0.0498s), its PSNR lags by 1.29 dB. Compared to DehazeFormer, the prior quality leader, our method reduces the inference time by 43.9% (0.0831s vs. 0.0498s) while improving the PSNR by 0.71 dB and using 53.7% fewer parameters. This dual optimization of accuracy and resource efficiency positioned MCAF-Net as the optimal choice for real-time haze removal in resource-constrained environments.
TABLE XI: Performance comparison of state-of-the-art methods
Method PSNR SSIM MSE #Time #Param FLOPs
FFA-Net 16.71 0.4792 0.0377 0.1200 4.456M 287.8G
GridDehazeNet 20.24 0.6312 0.0096 0.0699 955.7K 85.72G
4KDebazing 22.55 0.6912 0.0099 0.0549 34.55M 105.8G
SCAnet 19.07 0.5966 0.0180 0.0700 2.39M 258.6G
Trinity-Net 24.11 0.7103 0.0058 0.0400 20.14M 30.78G
DehazeFormer 24.69 0.7143 0.0051 0.0831 1.205M 39.76G
PCSformer 23.71 0.6547 0.0055 0.0600 3.73M 27.66G
PhDnet 24.28 0.6996 0.0053 0.0642 10.03M 33.24G
MCAF-Net 25.40 0.7221 0.0040 0.0498 558.1K 19.82G
VI Conclusion and Future Works
In this work, we present MCAF-Net, an innovative framework for real-world RSID, designed to overcome the limitations of synthetic-to-real domain gaps through a multi-branch adaptive architecture. Our approach integrates MFIBA, CSAM, and MFAFM to achieve comprehensive feature learning, precise color restoration, and context-aware haze suppression. The introduction of RRSHID dataset—the first large-scale real-world benchmark—provides critical support for modeling complex atmospheric variations and sensor-specific degradations. Extensive validation demonstrates that MCAF-Net not only outperforms existing methods in real-world scenarios but also maintains competitive efficiency with minimal computational overhead, offering practical value for time-sensitive earth observation tasks. While the current framework excels in static haze removal, challenges remain in handling dynamic atmospheric interactions and multi-modal sensor data. Future efforts will focus on integrating temporal modeling and cross-spectral fusion to advance toward all-weather, physics-aware RS restoration systems. Additionally, we recognize that the ultimate value of RSID lies in enhancing the performance of downstream tasks. The RRSHID dataset and MCAF-Net method have the potential to significantly improve applications such as object detection, change detection, and land cover classification under hazy conditions. Haze typically reduces the accuracy of these tasks, especially for small objects and low-contrast features. Our preliminary analysis indicates that using MCAF-Net as a preprocessing step can improve the performance of these tasks. In future work, we plan to construct benchmark datasets with specific task annotations to quantitatively evaluate the impact of dehazing on downstream applications, further validating the practical value of our proposed method in real-world RS applications.